ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 105, 1998

Abarbanel, J. On the Schur multiplier of augmented algebras  197

Alekseevsky, D., A. Kriegl, P. W. Michor and M. Losik Choosing roots of polynomials smoothly

203

Ara, P., K. R. Goodearl, K. C. O’Meara and E. Pardo Separative cancellation for projective modules over exchange rings

105

Campi, S., A. Colesanti and P. Gronchi Blaschke-decomposable convex bodies

185

Colesanti, A. See S. Campi

Corvaja, P. and U. Zannier Values of rational functions on nonhilbertian fields and a question of Weissauer

323

Deville, R., V. Fonf and P. Hájek Analytic and polyhedral approximation of convex bodies in separable polyhedral Banach spaces

139

Fonf, V. See R. Deville

Friedgut, E. and J. Kahn On the number of copies of one hypergraph in another

251

Goodearl, K. R. See P. Ara

Gronchi, P. See S. Campi

Guenais, M. Une majoration de la multiplicité spectrale d’opérateurs associés à des cocycles réguliers

263

Hájek, P. See R. Deville

Kahn, J. See E. Friedgut

Kenyon, R. Tilings and discrete Dirichlet problems  61
Kierstead, H. A. On-line coloring k-colorable graphs  93

Komorowski, R. A. and N. Tomczak-Jaegermann Erratum to: “Banach spaces without local unconditional structure”

85

Kriegl, A. See D. Alekseevsky

Lévai, L. Linear core conditions in residually finite groups  337

Levin, M. Certain finite dimensional maps and their application to hyperspaces

257

Losik, M. See D. Alekseevsky

Meiri, D. Entropy and uniform distribution of orbits in 𝕋^{d}  155

Michor, P. W. See D. Alekseevsky Nagy, P. T. and K. Strambach Loops, their cores and symmetric

spaces

285

Novo, S. and R. Obaya On the dynamical behaviour of an almost periodic linear system with an ergodic 2-sheet studied by R. Johnson

235

Obaya, R. See S. Novo O’Meara, K. C. See P. Ara Pardo, E. See P. Ara

Peretz, R. The geometry of the asymptotics of polynomial maps  1

Strambach, K. See P. T. Nagy Tomczak-Jaegermann, N. See R. A. Komorowski Zannier, U. See P. Corvaja